# Identifying non-Hermitian critical points with quantum metric

### v1

@@ Abstract
The geometric properties of quantum states is fully encoded by the quantum geometric tensor. The real and imaginary parts of the quantum geometric tensor are the quantum metric and Berry curvature, which characterize the distance and phase difference between two nearby quantum states in Hilbert space, respectively. For conventional Hermitian quantum systems, the quantum metric corresponds to the fidelity susceptibility and has already been used to specify quantum phase transitions from the geometric perspective. In this work, we extend this wisdom to the non-Hermitian systems for revealing non-Hermitian critical points. To be concrete, by employing numerical exact diagonalization and analytical methods, we calculate the quantum metric and corresponding order parameters in various non-Hermitian models, which include two non-Hermitian generalized Aubry-André models and non-Hermitian cluster and mixed-field Ising models. We demonstrate that the quantum metric of eigenstates in these non-Hermitian models exactly identifies the localization transitions, mobility edges, and many-body quantum phase transitions, respectively. We further show that this strategy is robust against the finite-size effect and different boundary conditions.
@@ I Introduction
Quantum phase transitions are a kind of phase transitions that are driven by quantum fluctuations and occur at zero temperature Sachdev (2011). They are extensively investigated between different magnetic phases Smacchia et al. (2011), localization phases Ganeshan et al. (2015); An et al. (2021); Wang et al. (2022); Tang et al. (2021), topological phases Matsumoto et al. (2020); Longhi (2019); Gong et al. (2018); Zhang et al. (2021), and so on. Near the phase transition point, the infinitesimal variation of certain system parameters will dramatically change the properties of the ground (eigen) state, which intuitively leads to a sharp change of the state fidelity. Moreover, an information-geometric approach based on this observation has already been put forward Zanardi and Paunković (2006); Zhu (2006, 2008); Zanardi et al. (2007a); You et al. (2007); Gu (2010); Ma et al. (2010); Carollo et al. (2020). The concrete procedure is to calculate the fidelity susceptibility, which defines the response of fidelity to the driving parameter of the system Hamiltonian near critical points  Zanardi and Paunković (2006); Zhu (2006, 2008); Zanardi et al. (2007a); You et al. (2007); Gu (2010).
@@
The fidelity susceptibility is closely related to the quantum metric, which characterizes the distance between two nearby quantum states in Hilbert space Kolodrubetz et al. (2017); Ding et al. (2024, 2022); Törmä (2023). The quantum metric is the real part of a more general geometric concept, called the quantum geometric tensor (QGT) Kolodrubetz et al. (2017); Provost and Vallee (1980), of which the imaginary part is the well-studied Berry curvature Thouless et al. (1982); Simon (1983); Berry (1984) for plenty of celebrated physical effects Guo et al. (2016); Sundaram and Niu (1999); Nagaosa et al. (2010); Aharonov and Bohm (1959); Hasan and Kane (2010); Qi and Zhang (2011); Zhang et al. (2018). The quantum metric is also related to many fascinating physical phenomena, such as the superfluidity in flat bands Julku et al. (2016); He et al. (2021) and topological quantum phases Roy (2014); Lim et al. (2015); Palumbo and Goldman (2018). The quantum metric has recently been experimentally measured in various artificial quantum systems Tan et al. (2019, 2021); Liao et al. (2021a); Chen et al. (2022); Asteria et al. (2019); Yu et al. (2022); Zheng et al. (2022); Lysne et al. (2023); Tan et al. (2018); Yu et al. (2019, 2024); Gianfrate et al. (2020); Yi et al. (2023); Cuerda et al. .
@@
Most of these previous studies focus on the QGT and related quantum phase transitions in Hermitian systems Solnyshkov et al. (2021); Campos Venuti and Zanardi (2007); Zanardi et al. (2007b); Zhang et al. ; Abasto et al. (2008); Albuquerque et al. (2010); Dey et al. (2012). In recent years, the theoretical and experimental explorations of non-Hermitian physics have garnered significant attentions Hatano and Nelson (1996, 1997); Bender and Boettcher (1998); Liu et al. (2019); Kawabata et al. (2019); Chen and Zhai (2018); Zhang et al. (2020a); Hu et al. ; Tu et al. (2023); Tzeng et al. (2021); Sticlet et al. (2023); Dóra et al. (2022); Yao and Wang (2018); Jiang et al. (2019); Song et al. (2019); Tang et al. (2020); Zhang et al. (2020b); Liu et al. (2020a); Zhai et al. (2020); Liu et al. (2020b); Hamazaki et al. (2019); Li et al. (2023); Liao et al. (2021b). Various novel phenomena or phases unique to non-Hermitian systems have been revealed, such as exceptional points Tzeng et al. (2021); Sticlet et al. (2023); Dóra et al. (2022), non-Hermitian skin effects Yao and Wang (2018); Song et al. (2019); Jiang et al. (2019), and non-Hermitian delocalization Tang et al. (2020); Zhang et al. (2020b); Liu et al. (2020a); Zhai et al. (2020); Liu et al. (2020b); Hamazaki et al. (2019). It is worth mentioning that phase transitions could be also sensitive to the non-Hermiticity, and the associated critical points in non-Hermitian systems have potential for sensitivity enhancement in the quantum sensing regime Li et al. (2023); Liao et al. (2021b); Di Candia et al. (2023); Liang et al. (2020); He et al. (2023).
@@
Here we exploit the information-geometric approach to study quantum phase transitions in both single-particle and many-body non-Hermitian systems. We explore the critical points in various non-Hermitian models through the quantum metric. By using the numerical exact diagonalization, we calculate the quantum metric and localization properties (the fractal dimension and participation ratio) in two non-Hermitian generalized Aubry-André (GAA) models Harper (1955); Aubry and André (1980). Our results show that the quantum metric will be divergent or singular in proximity of localization transition points for all eigenstates. Thus the quantum metric exactly identifies the localization transition and the mobility edges in these non-Hermitian quasiperiodic systems. For non-Hermitian cluster and mixed-field Ising models, we demonstrate that the quantum metric of the ground state can reveal the quantum phase transitions, as compared with the corresponding order parameters obtained from both analytical and numerical approaches. Finally, we show that this information-geometric strategy for identifying non-Hermitian critical points is robust against the finite-size effect and different boundary conditions.
@@ II Non-Hermitian quantum metric
We begin by considering a non-Hermitian Hamiltonian $H(\bm{\lambda})$ parameterized by $\bm{\lambda}=(\lambda_{1},\lambda_{2},\dots,\lambda_{D})$ in $D$-dimensional parameter space. The eigenstate equation is $H(\bm{\lambda})\ket{\psi_{n}^{R}(\bm{\lambda})}=E_{n}\ket{\psi_{n}^{R}(\bm{\lambda})}$ with $n$ the band index. The eigenvalues $E_{n}$ are in general complex due to the non-Hermiticity of the Hamiltonian, and the eigenstate |$\ket{\psi_{1}^{R}(\bm{\lambda})}$ corresponding to the minimum real eigenvalue is set as the ground state. The eigenstates satisfy the normalization condition $\braket{\psi_{n}^{R}(\bm{\lambda})}{\psi_{n}^{R}(\bm{\lambda})}=1$ but is generally nonorthogonal $\braket{\psi_{m}^{R}(\bm{\lambda})}{\psi_{n}^{R}(\bm{\lambda})}\neq 0$ $(m\neq n)$. We assume the ground state |$\ket{\psi_{1}^{R}(\bm{\lambda})}$ is non-degenerate and focus on the right eigenstates |$\ket{\psi_{n}^{R}(\bm{\lambda})}\doteq\ket{\psi_{n}(\bm{\lambda})}$ for simplicity. The distance between two neighbor eigenstates |$\ket{\psi_{n}(\bm{\lambda})}$ and |$\ket{\psi_{n}(\bm{\lambda}+d\bm{\lambda})}$ in the parameter space is given by Ye et al. ; Provost and Vallee (1980); Kolodrubetz et al. (2017)
@@
where $g^{(n)}_{\mu\nu}$ is the non-Hermitian quantum metric with parameters $\{\mu,\nu\}\in(\lambda_{1},\lambda_{2},\dots,\lambda_{D})$. It is the real part of the non-Hermitian quantum geometric tensor  Ye et al. ; Provost and Vallee (1980); Kolodrubetz et al. (2017)
@@
Here $F_{n}$ is the so-called fidelity of the $n$-th eigenstate, and $\chi^{(n)}_{F}$ is the fidelity susceptibility Zanardi and Paunković (2006); Zanardi et al. (2007a); Carollo et al. (2020). Notably, $\chi^{(n)}_{F}$ denotes the response of $F_{n}$ due to the small change of $\mu$. Combining Eq. (1) and Eq. (3), we can proof the equivalence of the quantum metric and the fidelity susceptibility in generic non-Hermitian Hamiltonians:
$\begin{split}g^{(n)}_{\mu\mu}&=\braket{\partial_{\mu}\psi_{n}}{\partial_{\mu}\psi_{n}}-\braket{\partial_{\mu}\psi_{n}}{\psi_{n}}\braket{\psi_{n}}{\partial_{\mu}\psi_{n}}\\ &=\lim_{d\mu\to 0}\frac{-2\ln F_{n}}{d\mu^{2}}=\chi^{(n)}_{F}.\end{split}$ (4)
It has been shown that the fidelity susceptibility of the ground state can detect quantum critical points in Hermitian systems  Zanardi and Paunković (2006); Zhu (2006, 2008); Zanardi et al. (2007a); You et al. (2007); Gu (2010). For $n$-th eigenstate, one can also expect that the corresponding quantum metric $g^{(n)}_{\mu\mu}$ will reach the maximum value or diverge at its localization transition point. In the following, we extend this wisdom to generic non-Hermitian models. To be concrete, we demonstrate that the localization transition of all eigenstates with related mobility edges and the quantum phase transitions in non-Hermitian disordered and many-body systems can be revealed by the quantum metric, respectively.
@@ III.1 Detecting localization transitions
with $V_{1}$ and $V_{2}$ being the modulation amplitudes of the potential and hopping terms, respectively, $h$ being the complex phase, and $\beta$ being an irrational number to ensure the incommensurate modulation. Hereafter, we choose $\beta=(\sqrt{5}-1)/2$ as the golden ratio and the lattice site number $L$ from the Fibonacci sequence, and set $t=1$ as the unit of energy.
@@
$\eta_{n}=-\frac{\ln(\sum_{j}^{L}|\braket{j}{\psi_{n}}|^{2})}{\ln L},$ (7)
where |$\ket{j}$ is the computational basis and |$\ket{\psi_{n}}$ is the $n$-th right eigenstates. In this model, there is no mobility edge and all eigenstates share the same localization properties Chang et al. (1997); Liu et al. (2015); Wang et al. (2021); Tang et al. (2021). Thus, we focus on the ground state |$\ket{\psi_{1}}$ with the corresponding fractal dimension $\eta_{1}\equiv\eta$ for simplicity. When the system size $L\to\infty$, one has $\eta\to 1$ for the extended phase, $\eta\to 0$ for the localized phase, while $\eta$ falls between 00 and $1$ ($\eta\approx 0.5$) in the critical phase, respectively. Figure 1(a) show the numerical results of $\eta$ in the parameter space ($V_{1}$,$V_{2}$) for nonreciprocal case with $g=0.5$ and $h=0$ under the periodic boundary condition (PBC), by using the exact diagonalization method. Three regimes (denoted by the I, II and III) for the extended, critical, and localized phases are exhibited. Notably, the critical values $V_{1c}$ for localization transition points with respect to other parameters are given by Tang et al. (2021)
@@
To demonstrate the ability of the non-Hermitian quantum metric to detect localization transition points in non-Hermitian cases, we numerically calculate the diagonal elements of the quantum metric $g_{\mu\mu}$ for the parameter $\mu=\{V_{1},V_{2}\}$. To resole the three different localization regions in this model, we use the logarithm quantity of the quantum metric

### v3

@@ Abstract
The geometric properties of quantum states is fully encoded by the quantum geometric tensor. The real and imaginary parts of the quantum geometric tensor are the quantum metric and Berry curvature, which characterize the distance and phase difference between two nearby quantum states in Hilbert space, respectively. For conventional Hermitian quantum systems, the quantum metric corresponds to the fidelity susceptibility and has already been used to specify quantum phase transitions from the geometric perspective. In this work, we extend this wisdom to the non-Hermitian systems for revealing non-Hermitian critical points. To be concrete, by employing numerical exact diagonalization and analytical methods, we calculate the quantum metric and corresponding order parameters in various non-Hermitian models, which include two non-Hermitian generalized Aubry-André models and non-Hermitian cluster and mixed-field Ising models. We demonstrate that the quantum metric of eigenstates in these non-Hermitian models exactly identifies the localization transitions, mobility edges, and many-body quantum phase transitions with gap closings, respectively. We further show that this strategy is robust against the finite-size effect and different boundary conditions.
@@ I Introduction
Quantum phase transitions are a kind of phase transitions that are driven by quantum fluctuations and occur at zero temperature Sachdev (2011). They are extensively investigated between different magnetic phases Smacchia et al. (2011), localization phases Ganeshan et al. (2015); An et al. (2021); Wang et al. (2022); Tang et al. (2021), topological phases Matsumoto et al. (2020); Longhi (2019); Gong et al. (2018); Zhang et al. (2021), and so on. Near the phase transition point, the infinitesimal variation of certain system parameters will dramatically change the properties of the ground (eigen) state, which intuitively leads to a sharp change of the state fidelity. Moreover, an information-geometric approach based on this observation has already been put forward Zanardi and Paunković (2006); Zhu (2006, 2008); Zanardi et al. (2007a); You et al. (2007); Gu (2010); Lv et al. (2022a, b); Ma et al. (2010); Carollo et al. (2020). The concrete procedure is to calculate the fidelity susceptibility originally defined in Ref. You et al. (2007), which characterizes the response of fidelity to the driving parameter of the system Hamiltonian near critical points.
@@
The fidelity susceptibility is closely related to the quantum metric, which characterizes the distance between two nearby quantum states in Hilbert space Kolodrubetz et al. (2017); Ding et al. (2024, 2022); Törmä (2023). The quantum metric is the real part of a more general geometric concept, called the quantum geometric tensor (QGT) Kolodrubetz et al. (2017); Provost and Vallee (1980), of which the imaginary part is the well-studied Berry curvature Thouless et al. (1982); Simon (1983); Berry (1984) for plenty of celebrated physical effects Guo et al. (2016); Sundaram and Niu (1999); Nagaosa et al. (2010); Aharonov and Bohm (1959); Hasan and Kane (2010); Qi and Zhang (2011); Zhang et al. (2018). The quantum metric is also related to many fascinating physical phenomena, such as the superfluidity in flat bands Julku et al. (2016); He et al. (2021) and topological quantum phases Roy (2014); Lim et al. (2015); Palumbo and Goldman (2018). The quantum metric has recently been experimentally measured in various artificial quantum systems Tan et al. (2019, 2021); Liao et al. (2021); Chen et al. (2022); Asteria et al. (2019); Yu et al. (2022); Zheng et al. (2022); Lysne et al. (2023); Tan et al. (2018); Yu et al. (2019, 2024); Gianfrate et al. (2020); Yi et al. (2023); Cuerda et al. .
@@
Most of these previous studies focus on the QGT and related quantum phase transitions in Hermitian systems Solnyshkov et al. (2021); Campos Venuti and Zanardi (2007); Zanardi et al. (2007b); Zhang et al. (a); Abasto et al. (2008); Hetényi and Lévay (2023); Albuquerque et al. (2010); Dey et al. (2012); Zhang et al. (b). In recent years, the theoretical and experimental explorations of non-Hermitian physics have garnered significant attentions Ashida et al. (2020); Hatano and Nelson (1996, 1997); Bender and Boettcher (1998); Liu et al. (2019); Kawabata et al. (2019); Chen and Zhai (2018); Zhang et al. (2020a); Hu et al. ; Tu et al. (2023); Tzeng et al. (2021); Sticlet et al. (2023); Dóra et al. (2022); Yao and Wang (2018); Jiang et al. (2019); Song et al. (2019); Tang et al. (2020); Zhang et al. (2020b); Liu et al. (2020a); Zhai et al. (2020); Liu et al. (2020b); Hamazaki et al. (2019); Li et al. (2023); Liao et al. (2021); Li et al. (2024). Various novel phenomena or phases unique to non-Hermitian systems have been revealed, such as exceptional points Bergholtz et al. (2021); Tzeng et al. (2021); Sticlet et al. (2023); Dóra et al. (2022); Alvarez et al. (2018); Agarwal et al. (a, b), non-Hermitian skin effects Yao and Wang (2018); Song et al. (2019); Jiang et al. (2019), and non-Hermitian delocalization Tang et al. (2020); Zhang et al. (2020b); Liu et al. (2020a); Zhai et al. (2020); Liu et al. (2020b); Hamazaki et al. (2019). The non-Hermitian quantum metric can characterize the band geometry in pseudo-Hermitian systems that differs from its Hermitian counterpart Zhu et al. (2021). It also serves as a tool to detect the symmetry breaking in PT-symmetric systems Zhang et al. (2019) and becomes a crucial quantity in the dynamics of wave packets near the exceptional points of non-Hermitian systems, diverging in a manner that fully controls the trajectory description of the wave packets Solnyshkov et al. (2021). Recent studies have further shown that the quantum metric can characterize topological transitions in the non-Hermitian SSH model Li et al. (2023); Chen Ye et al. (2024), and the critical scaling near exceptional points can be harnessed to enhance sensing precision Li et al. (2023); Liao et al. (2021); Di Candia et al. (2023); Liang et al. (2020); He et al. (2023).
@@
Here we exploit the information-geometric approach to study quantum phase transitions in both single-particle and many-body non-Hermitian systems. We explore the critical points in various non-Hermitian models through the quantum metric. By using the numerical exact diagonalization, we calculate the quantum metric and localization properties (the fractal dimension and participation ratio) in two non-Hermitian generalized Aubry-André (GAA) models Harper (1955); Aubry and André (1980); Li et al. (2024). Our results show that the quantum metric will be divergent or singular in proximity of localization transition points for all eigenstates. Thus the quantum metric exactly identifies the localization transition and the mobility edges in these non-Hermitian quasiperiodic systems. For the non-Hermitian cluster and mixed-field Ising models, we demonstrate that the quantum metric of the ground state can reveal the quantum phase transitions with gap closing, as compared with the corresponding order parameters obtained from both analytical and numerical approaches. Finally, we show that this information-geometric strategy for identifying non-Hermitian critical points is robust against the finite-size effect and different boundary conditions.
@@ II Non-Hermitian quantum metric
We begin by considering a non-Hermitian Hamiltonian $H(\bm{\lambda})$ parameterized by $\bm{\lambda}=(\lambda_{1},\lambda_{2},\dots,\lambda_{D})$ in $D$-dimensional parameter space. The eigenstate equation is $H(\bm{\lambda})\ket{\psi_{n}^{R}(\bm{\lambda})}=E_{n}\ket{\psi_{n}^{R}(\bm{\lambda})}$ with $n$ the band index. The eigenvalues $E_{n}$ are in general complex due to the non-Hermiticity of the Hamiltonian, and the eigenstate |$\ket{\psi_{1}^{R}(\bm{\lambda})}$ corresponding to the minimum real eigenvalue is set as the ground state. The eigenstates satisfy the normalization condition $\braket{\psi_{n}^{R}(\bm{\lambda})}{\psi_{n}^{R}(\bm{\lambda})}=1$ but is generally nonorthogonal $\braket{\psi_{m}^{R}(\bm{\lambda})}{\psi_{n}^{R}(\bm{\lambda})}\neq 0$ $(m\neq n)$. We assume the ground state |$\ket{\psi_{1}^{R}(\bm{\lambda})}$ is non-degenerate and focus on the right eigenstates |$\ket{\psi_{n}^{R}(\bm{\lambda})}\doteq\ket{\psi_{n}(\bm{\lambda})}$ for simplicity. The distance between two neighbor eigenstates |$\ket{\psi_{n}(\bm{\lambda})}$ and |$\ket{\psi_{n}(\bm{\lambda}+d\bm{\lambda})}$ in the parameter space is given by Ye et al. ; Provost and Vallee (1980); Kolodrubetz et al. (2017); Matsumoto et al. (2020); Sun et al. (2022)
@@
where $g^{(n)}_{\mu\nu}$ is the non-Hermitian quantum metric with parameters $\{\mu,\nu\}\in(\lambda_{1},\lambda_{2},\dots,\lambda_{D})$. It is the real part of the non-Hermitian quantum geometric tensor Matsumoto et al. (2020); Sun et al. (2022)
@@
Here $F_{n}$ is the so-called fidelity of the $n$-th eigenstate, and $\chi^{(n)}_{F}$ is the fidelity susceptibility You et al. (2007). The fidelity susceptibility $\chi^{(n)}_{F}$ denotes the response of $F_{n}$ due to the small change of $\mu$. For ground states in Hermitian systems, the equivalence of the fidelity susceptibility and the quantum metric element has been shown. For continuous quantum phase transitions with gap closing, the fidelity susceptibility and quantum metric can be used to detect the critical points Zanardi and Paunković (2006); Zhu (2006, 2008); Zanardi et al. (2007a); You et al. (2007); Lv et al. (2022a); Gu (2010); Lv et al. (2022b). However, this wisdom may fail for discontinuous transitions without gap closing. For instance, it was found that the ground-state fidelity susceptibility is unable to witness the extended-critical phase transition point in the Hermitian p-wave-pair Aubry-André-Harper model due to the absence of gap closing Lv et al. (2022b).
Combining Eq. (1) and Eq. (3) for self-normal right eigenstates, we can find the equivalence of the diagonal element of quantum metric and the fidelity susceptibility in generic non-Hermitian Hamiltonians:
$\begin{split}g^{(n)}_{\mu\mu}&=\text{Re}(\braket{\partial_{\mu}\psi_{n}}{\partial_{\mu}\psi_{n}}-\braket{\partial_{\mu}\psi_{n}}{\psi_{n}}\braket{\psi_{n}}{\partial_{\mu}\psi_{n}})\\ &=\lim_{d\mu\to 0}\frac{-2\ln F_{n}}{d\mu^{2}}=\chi^{(n)}_{F}.\end{split}$ (4)
For $n$-th eigenstate, one can expect that the corresponding quantum metric $g^{(n)}_{\mu\mu}$ will reach the maximum value or diverge at its localization transition point. In the following, we extend this wisdom to generic non-Hermitian single particle quasi-periodic models and many-body spin models. To be concrete, we demonstrate that the localization transition of all eigenstates with related mobility edges and the quantum phase transitions (with gap closing) in non-Hermitian disordered and many-body systems can be revealed by the quantum metric, respectively.
@@ III.1 Detecting localization transitions
with $V_{1}$ and $V_{2}$ being the modulation amplitudes of the potential and hopping terms, respectively, $h$ being the complex phase, and $\beta$ being an irrational number to ensure the incommensurate modulation. Without loss of generality, we choose $\beta=(\sqrt{5}-1)/2$ as the golden ratio, which can be obtained from the Fibonacci sequence $F_{j+1}=F_{j}+F_{j-1}$ as $\beta=\lim_{j\to\infty}F_{j}/F_{j+1}$. In numerical simulations, we set the lattice site $L=610$ from $F_{j}$ to satisfy the periodic boundary condition (PBC), which is large enough to detect the critical points for $H_{\text{GAA}}^{(1)}$. This will be clearly seen from the finite-size scaling in Fig. 7 in Sec. V. We set $t=1$ as the unit of energy.
@@
$\eta_{n}=-\frac{\ln(\sum_{j}^{L}|\braket{j}{\psi_{n}}|^{4})}{\ln L},$ (7)
where |$\ket{j}$ is the computational basis and |$\ket{\psi_{n}}$ is the $n$-th right eigenstates. In this model, there is no mobility edge and all eigenstates share the same localization properties Chang et al. (1997); Liu et al. (2015); Wang et al. (2021); Tang et al. (2021). Thus, we focus on the ground state |$\ket{\psi_{1}}$ with the corresponding fractal dimension $\eta_{1}\equiv\eta$ for simplicity. When the system size $L\to\infty$, one has $\eta\to 1$ for the extended phase, $\eta\to 0$ for the localized phase, while $\eta$ falls between 00 and $1$ ($\eta\approx 0.5$) in the critical phase, respectively. Figure 1(a) show the numerical results of $\eta$ in the parameter space ($V_{1}$,$V_{2}$) for nonreciprocal case with $g=0.5$ and $h=0$ under the PBC, by using the exact diagonalization method. Three regimes (denoted by the I, II and III) for the extended, critical, and localized phases are exhibited. Notably, the critical values $V_{1c}$ for localization transition points with respect to other parameters are given by Tang et al. (2021)
@@
To demonstrate the ability of the non-Hermitian quantum metric to detect localization transition points in non-Hermitian cases,we numerically calculate the diagonal elements of the quantum metric $g_{\mu\mu}$ of ground state for the parameter $\mu=\{V_{1},V_{2}\}$, which are equivalent to the ground-state fidelity susceptibility. To resolve the three different localization regions in this model, we use the logarithm quantity of the quantum metric